# Understanding Software Engineering Agents: A Study of Thought-Action-Result Trajectories

### v1

@@ III-C RQ3: Semantic Relationships
RepairAgent 98.6 1.4 1.4 6.3 69.7 9.2 13.5 17.8 51.4 20.2 10.6 75.4 3.4 3.4 17.8 30.0 5.6 64.4
- Success 99.2 0.8 1.4 4.2 75.1 5.2 14.2 26.1 56.7 11.8 5.3 82.5 4.0 1.7 11.8 43.5 0.9 55.6
- Failure 98.6 1.4 1.4 6.6 68.9 9.8 13.4 16.7 50.7 21.3 11.3 74.5 3.3 3.6 18.6 27.4 6.5 66.1
OpenHands 96.8 3.2 0.1 8.1 80.2 1.6 10.1 23.8 61.2 13.7 1.2 70.6 3.3 6.8 19.3 30.3 5.1 64.6
- Success 99.0 1.0 0.0 5.9 78.6 1.8 13.6 28.6 58.7 12.2 0.5 73.4 2.8 2.8 21.0 19.6 2.7 77.8
- Failure 96.5 3.5 0.1 8.3 80.4 1.6 9.7 23.3 61.6 13.9 1.3 70.3 3.3 7.4 19.0 31.4 5.3 63.3
ACR 96.3 3.7 0.0 1.4 72.4 0.9 25.3 20.8 69.7 3.9 5.6 59.9 0.0 6.2 33.9 37.7 25.8 36.5
- Success 100.0 0.0 0.0 2.1 70.8 0.0 27.1 19.5 78.0 2.4 0.0 59.7 0.0 8.1 32.3 46.7 21.7 31.7
- Failure 95.2 4.8 0.0 1.2 72.8 1.2 24.9 21.2 67.2 4.4 7.3 59.9 0.0 5.7 34.4 34.8 27.2 38.0
Table II presents the distribution of semantic relationships across trajectories for the three agents. Columns represent key relational categories (e.g., Divergence, Follow-up, Refinement…). Rows show the prevalence of each relation for both successful and unsuccessful trajectories.
@@ III-C1 Analysis of Thought-Action Relationships
Among the three agents, misalignment between thought and action is rare but closely associated with failing trajectories. For example, OpenHands shows 1% misalignment in successful and 3.5% in failing trajectories, while AutoCodeRover has 0% in successful and 4.8% in failing ones. Although not all misalignments result in failure, even a single instance can significantly increase trajectory length or cause failure. Furthermore, misalignment is more prevalent in failing trajectories: only 1 out of 10 successful RepairAgent trajectories contain a misalignment compared to 12 out of 30 failing ones.
An example of misalignment between intent and action occurs at iteration 5 during RepairAgent’s execution on bug Compress_13. The agent, after identifying a plausible fix, attempted refinement but instead suggested an empty fix. This error caused confusion and extended the process from iteration 6 (where the correct patch was available) to iteration 38, increasing computational cost. Although such misalignment is uncommon and model-dependent, incorporating mechanisms to verify that actions align with articulated thoughts is prudent. Prior work has explored this for LLMs and agents [38, 39]. Newer reasoning models (e.g., o3, deepseek, Claude 4) embed self-reflection and criticism. Integrating these into SE agents, where precise translation of natural language to actions is critical, would improve robustness and efficiency in a significant fraction of trajectories.
Thought-Action Alignment Even a single misalignment between thought and action can cause failure or increase computational cost. We recommend explicit verification alignment, e.g, through self-reflection, critique frameworks, or reasoning models.
@@ III-C2 Analysis of Thought-Thought Relationships
Unlike chain-of-thought reasoning, LLM-agent thoughts are interleaved with actions and their outcomes, influencing subsequent thoughts. Table II shows that most thoughts are follow-ups or refinements of previous ones, though distinct patterns emerge per agent. For example, RepairAgent has the highest number of contradicting thoughts, which are rare in OpenHands and AutoCodeRover. GPT-3.5, used by RepairAgent, often makes trivial mistakes. In bug Compress_13, RepairAgent states at iteration 5: The fix was successful, and all test cases passed…, but contradicts this at iteration 6: The previous fixes did not work, and the bug is still present…. Additionally, RepairAgent frequently produces redundant thoughts, especially in failing trajectories, often repeating generic statements like The previous fix failed, I need to suggest a new one without explaining failure reasons or next steps.
Divergence between thoughts is notable in RepairAgent (37/40 trajectories) and OpenHands (34/40), with both positive and negative effects. Diverging from promising paths may mislead the agent, while diverging from unproductive loops enables exploration and balances exploitation and exploration. For example, in OpenHands’s Django-10924, iterations 1–6 focus on FilePathField implementation, but iteration 7 shifts to reproducing the issue, breaking repetitive code reading and exploring testing and fixing instead.
@@ III-C3 Analysis of Consecutive Actions Pairs
Across all three agents, Follow-up relationships dominate, indicating most actions build on their immediate predecessors. However, the distribution of other relationship types varies among agents. Divergence is prominent in Action-Action relationships and is often beneficial, reflecting transitions between phases (e.g., from bug localization to fix suggestion). This is evident as action divergence is more common in successful trajectories (e.g., 26 vs. 16 for RepairAgent). Conversely, repetitive sequences and consecutive action repetition, particularly in RepairAgent and AutoCodeRover, strongly correlate with failing trajectories. Repetition wastes iterations and may induce unproductive loops, as also observed in RQ2. Repetition can be detected using sequence analysis methods such as sliding window matching or n-gram frequency analysis. Detected repetitions can then be addressed by diversification strategies that encourage exploration of alternative actions or transitions.
@@ III-C4 Influence of Results on Thoughts and Actions
Across all agents, the most common relationship is follow-up, where a thought logically builds on the received result, such as responding to a failed test by suggesting a new patch. Refinement relationships are also frequent, indicating efforts to improve the solution based on results. Misinterpretations and no influence cases are less common but occur more often in failing trajectories, reflecting instances where agents ignore or misunderstand result signals. For example, in RepairAgent’s run on bug Cli_11 at iteration 6, the agent misinterprets the result indicating that the last fix was empty and instead of suggesting a non-empty fix, it moves on to analyzing the code with the intention to change empty function arguments.
@@
However, failures often correlate with a higher proportion of no-influence actions (0.9 vs 6.5 for RepairAgent), where the agent’s next move does not reflect the information content of the preceding result. These observations suggest that effective LLM agents not only need to interpret results accurately at the reasoning level but also translate those insights into appropriately responsive action sequences.
@@ IV Limitations and Threats to Validity
Our study offers insights into LLM-based agents in software engineering tasks but faces limitations, notably the potential subjectivity in our annotation and classification process despite inter-annotator agreement checks. We evaluate three bug-fixing agents which, despite their common use, do not fully capture all software development tasks or real-world deployment scenarios. The agents rely on distinct models, ensuring diversity but complicating fair comparisons. The generalization of our results to other SE agents or tasks, such as code generation or testing, remains open for study.
@@ VI Conclusion
This study systematically analyzes LLM-based autonomous agents in software engineering, focusing on decision-making, reasoning consistency, and trajectory dynamics. Using statistical analysis, semantic relation modeling, open coding, and sequence pattern mining, we identify key behavioral trends. Successful trajectories balance information gathering, hypothesis testing, and fix validation, while unsuccessful ones exhibit redundant exploration or premature fixes. Categorizing actions and sequences reveals agents’ strengths and weaknesses, guiding improvements in efficiency, reliability, and reasoning. Future work could include expanding agent and benchmark coverage, refining classification methods, and developing automatic failure detection and mitigation techniques.
@@
The paper is under review. We will release the data once our work is published in its final version.

### v2

@@ III-C RQ3: Semantic Relationships
RepairAgent 98.5 1.5 1.4 6.3 63.8 13.3 15.1 20.6 50.9 16.2 12.3 76.1 3.6 3.9 16.4 25.8 5.3 68.9
- Success 97.7 2.3 1.2 4.1 73.1 6.9 14.7 23.3 57.9 12.7 6.1 81.5 4.8 2.0 11.7 41.2 1.2 57.6
- Failure 98.7 1.3 1.4 6.8 61.9 14.7 15.2 20.0 49.4 17.0 13.6 74.9 3.4 4.3 17.4 22.5 6.2 71.3
OpenHands 98.7 1.3 0.0 14.8 72.8 1.9 10.5 25.5 57.0 14.2 3.2 72.1 3.9 6.3 17.6 29.2 3.9 66.9
- Success 99.5 0.5 0.0 8.6 74.5 2.0 14.8 28.6 55.6 15.3 0.5 73.9 2.3 6.1 17.7 29.6 2.8 67.6
- Failure 98.6 1.4 0.1 16.2 72.4 1.8 9.6 24.9 57.4 14.0 3.8 71.6 4.4 6.4 17.6 29.1 4.2 66.7
ACR 96.3 3.7 0.0 1.1 75.8 2.8 20.2 24.2 64.6 4.5 6.7 62.2 0.0 9.4 28.3 38.8 12.4 48.9
- Success 100.0 0.0 0.0 2.4 78.1 0.0 19.5 26.8 68.3 4.8 0.0 64.3 0.0 11.9 23.8 43.9 4.9 51.2
- Failure 95.2 4.8 0.0 0.7 75.2 3.7 20.4 23.4 63.5 4.4 8.8 61.6 0.0 8.7 29.7 37.2 14.6 48.2
Table II presents the distribution of semantic relationships across trajectories of the three agents. Columns represent key relational categories (e.g., Divergence, Follow-up, Refinement…). Rows show the prevalence of each relation for both successful and unsuccessful trajectories.
@@ III-C1 Analysis of Thought-Action Relationships
Among the three agents, misalignment between thought and action is rare but closely associated with failing trajectories. For example, OpenHands shows 0.5% misalignment in successful and 1.4% in failing trajectories, while AutoCodeRover has 0% in successful and 4.8% in failing ones. Although not all misalignments result in failure, even a single instance can significantly increase trajectory length or cause failure. Furthermore, misalignment is more prevalent in failing trajectories with 2.7 misalignment counts in failing trajectories vs one misalignment in successful trajectories of OpenHands (that contain at least one misalignment).
An example of misalignment between intent and action occurs at iteration 5 during RepairAgent’s execution on bug Compress_13. The agent, after identifying a plausible fix, attempted refinement but instead suggested an empty fix. This error caused confusion and extended the process from iteration 6 (where the correct patch was available) to iteration 38, increasing computational cost. Although such misalignment is uncommon and model-dependent, incorporating mechanisms to verify that actions align with articulated thoughts is prudent Prior work has explored this for LLMs and agents [38, 39]. Newer reasoning models (e.g., o3, deepseek, Claude 4) embed self-reflection and criticism. Integrating these into software engineering agents, where precise translation of natural language to actions is critical, would improve robustness and efficiency in a significant fraction of trajectories.
Thought-Action Alignment Even a single misalignment between thought and action can cause failure or increase computational cost. We recommend explicit validation that actions align with thoughts, e.g., through self-reflection, critique frameworks, or reasoning models.
@@ III-C2 Analysis of Thought-Thought Relationships
Unlike chain-of-thought reasoning, LLM-agent thoughts are interleaved with actions and their outcomes, influencing subsequent thoughts. Table II shows that most thoughts are follow-ups or refinements of previous ones, though distinct patterns emerge per agent. For example, RepairAgent has the highest number of contradicting thoughts, which are rare in OpenHands and AutoCodeRover. GPT-3.5, used by RepairAgent, often makes trivial mistakes. In bug Compress_13, RepairAgent states at iteration 5: The fix was successful, and all test cases passed., but contradicts this at iteration 6: The previous fixes did not work, and the bug is still present. Additionally, RepairAgent frequently produces redundant thoughts, especially in failing trajectories, often repeating generic statements like The previous fix failed, I need to suggest a new one without explaining failure reasons or next steps.
Divergence between thoughts is notable in RepairAgent (29/40 trajectories) and OpenHands (24/40), with both positive and negative effects. Diverging from promising paths may mislead the agent, while diverging from unproductive loops enables exploration and balances exploitation and exploration. For example, in OpenHands’s Django-10924, iterations 1–6 focus on understanding the implementation of FilePathField, but iteration 7 shifts to reproducing the issue, stopping repetitive code reading in favor of testing and fixing.
@@ III-C3 Analysis of Consecutive Actions Pairs
The actions taken by an agent should logically follow from prior actions, reflecting a coherent strategy. Across all three agents, Follow-up relationships dominate, indicating most actions build on their immediate predecessors. However, the distribution of other relationship types varies among agents. Divergence is prominent in Action-Action relationships and is often beneficial, reflecting transitions between phases (e.g., from bug localization to fix suggestion). This is evident as action divergence is higher in successful trajectories. Conversely, repetitive sequences and consecutive action repetition, particularly in RepairAgent and AutoCodeRover, strongly correlate with failing trajectories. Repetition wastes iterations and may induce unproductive loops, as also observed in RQ2. As shown in the table, action repetition is several times higher in failing trajectories. Repetition can be detected using sequence analysis methods such as sliding window matching or n-gram frequency analysis. Detected repetitions can then be addressed by diversification strategies that encourage exploration of alternative actions or transitions.
@@ III-C4 Influence of Results on Thoughts and Actions
Across all agents, the most common relationship is follow-up, where a thought logically builds on the received result, such as responding to a failed test by suggesting a new patch. Refinement relationships are also frequent, indicating efforts to improve the solution based on results. Misinterpretations and no influence cases are less common but occur in both failing and successful instances. For example, in RepairAgent’s run on bug Cli_11 at iteration 6, the agent misinterprets the result indicating that the last fix was empty and instead of suggesting a non-empty fix, it moves on to analyzing the code with the intention to change empty function arguments.
@@
However, failures often correlate with a higher proportion of no-influence actions (1.2% vs. 6.2% for RepairAgent and 4.9% vs 14.6% for AutoCodeRover), where the agent’s next move does not reflect the information content of the preceding result. These observations suggest that effective LLM agents not only need to interpret results accurately at the reasoning level but also translate those insights into appropriately responsive action sequences.
Result Driven Reasoning and Actions Improving LLM-agents requires mechanisms that enhance result sensitivity, ensuring results consistently guide both reasoning and action decisions towards suggesting an effective next actions.
@@ IV Limitations and Threats to Validity
Our study offers insights into LLM-based agents in software engineering tasks but faces limitations, notably the potential subjectivity in our annotation and classification process despite partial inter-annotator agreement checks. We evaluate three bug-fixing agents which, despite their common use, do not fully capture all software development tasks or real-world deployment scenarios. The agents rely on distinct models, ensuring diversity but complicating fair comparisons. The generalization of our results to other software engineering agents or tasks, such as code generation or testing, remains open for study.
@@ VI Conclusion
This study systematically analyzes LLM-based autonomous agents in software engineering, focusing on decision-making, reasoning consistency, and trajectory dynamics. Using statistical analysis, semantic relation modeling, open coding, and sequence pattern mining, we identify key behavioral trends. Successful trajectories balance information gathering, hypothesis testing, and fix validation, while unsuccessful ones exhibit redundant exploration or premature fixes. Categorizing actions and sequences reveals strengths and weaknesses of agents, guiding improvements in efficiency, reliability, and reasoning. Future work could include expanding agent and benchmark coverage, refining classification methods, and developing automatic failure detection and mitigation techniques.
@@
Our code and data are available at:
https://github.com/sola-st/llm-agents-study
Acknowledgments
This work was supported by the European Research Council (ERC, grant agreements 851895 and 101155832) and by the German Research Foundation within the DeMoCo project.